# Reward-Safety Balance in Offline Safe RL via Diffusion Regularization

### v1

@@ 4 Practical Algorithm
In this section, we present the detailed procedure for the implementation of our main algorithm Diffusion-Regularized Constrained Offline Safe Reinforcement Learning (DRCORL), as outlined in Algorithm 1, where we provide two candidates for the SafeAdaptation step outlined in Algorithms 2 and 3 given in Appendix B. This includes both the pretraining stage and the policy extraction stage.
@@ 4.1 Pretraining the Diffusion Policy and Critics
$\min_{\pi\in\Pi}\mathbb{E}_{{s},{a},{s},a^{\prime}\sim\pi(\cdot|s^{\prime})}\big{[}\big{\|}r({s},{a})+\gamma Q^{\pi}_{c}(s^{\prime},a^{\prime})-Q^{\pi}_{c}(s,a)\|^{2}\big{]}-\alpha\mathbb{E}_{s,a\sim\pi}[Q^{\pi}_{c}(s,a)].$ (19)
@@ 4.2 Extracting Policy
An alternative to the cost optimization objective Eq. (21) is to incorporate the primal-dual based objective into our optimization process. By adding Lagrangian multiplier $\lambda$ into Eq. (21), we can obtain the new objective
$\max\mathbb{E}_{s,a\sim\pi_{\theta}}\left[Q^{\pi_{\theta}}_{r}(s,a)-\lambda(Q^{\pi_{\theta}}_{c}(s,a)-l)-\frac{1}{\beta}l(\pi_{\theta},\mu_{\psi}|s)\right].$ (22)
@@
We can update this multiplier $\lambda$ using the PID method Stooke et al. (2020). The advantage of leveraging the multiplier is that we can align the reward objective with the cost objective to avoid instability during the training. Nevertheless, in this work we first consider the case under the cost optimization objective in Eq. (21). Using the result obtained in Section 3.2, we can obtain the gradient of Eq. (20) and Eq. (21) with respect to $\theta$ using the score function of the diffusion model $\epsilon_{\psi}$ and the critic function. We take the Gaussian policy family under a constant variance. For example, with $\Pi:=\{\pi_{\theta}(a|s)=\mathcal{N}(a;m_{\theta}(s),\Sigma_{\theta}(s)\}$, we can simplify the gradient to
$g_{r}$ $=\mathbb{E}_{{s},{a}\sim\pi_{\theta}}\left[\big{(}\nabla_{{a}}Q^{\pi_{\theta}}_{r}({s},{a})+\frac{1}{\beta}h_{\psi}(s,a)\big{)}\nabla_{\theta}\pi_{\theta}({s})\right],$ (23)
$g_{c}$ $=\mathbb{E}_{{s},{a}\sim\pi_{\theta}}\left[\big{(}-\nabla_{{a}}Q^{\pi_{\theta}}_{c}({s},{a})+\frac{1}{\beta}h_{\psi}(s,a)\big{)}\nabla_{\theta}\pi_{\theta}({s})\right],$
where $h_{\psi}(s,a)=\frac{1}{\sqrt{\bar{\beta}_{t}}}\epsilon_{\psi}(a_{t},t|s)|_{t\rightarrow 0}$. In Eq. (23), $a_{t}$ denotes the action with perturbed noise, $a_{t}=\sqrt{\bar{\alpha}_{t}}a+\sqrt{\bar{\beta}_{t}}\epsilon_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,I)$, and the notation $\nabla_{\theta}\pi_{\theta}(s)$ denotes the gradient of the action given state $s$ with respect to $\theta$. By assuming $a=m_{\theta}(s)+\Sigma^{1/2}_{\theta}(s)z$ where $z\sim\mathcal{N}(0,I)$, we obtain that $\nabla_{\theta}\pi_{\theta}(s)=\nabla_{\theta}m_{\theta}(s)+\nabla_{\theta}\Sigma^{1/2}_{\theta}(s)z$.
Our problem involves two optimization objectives: on one hand, we aim to maximize the reward as stated in Eq. (10); on the other hand, we seek to minimize the cost when the safety condition is violated, as described in Eq. (21). To address this, we consider two potential approaches. The first approach builds on the idea of CRPO Xu et al. (2021) and is specified in Algorithm 2, which switches between the two optimization objectives. Specifically, at each step, we evaluate the cost value $V_{c}^{\pi}(\rho)$. If the cost value does not exceed the threshold $l+h$, we continue optimizing the reward. Otherwise, we switch to optimizing the cost. Here, $h$ acts as a slack variable, serving as a soft boundary between reward and cost optimization. The second approach utilizes the gradient manipulation technique outlined in Gu et al. (2024a) and is specified in Algorithm 3, which was originally applied for safe adaptation in an online safe RL setting. In this method, we introduce two slack variables, $h^{-}$ and $h^{+}$. If the cost value satisfies $V_{c}^{\pi}(\rho)\leq l-h^{-}$, the policy is considered to be in a completely safe region, and we focus solely on maximizing the reward. Conversely, if the cost value satisfies $V_{c}^{\pi}(\rho)\geq l+h^{+}$, we prioritize minimizing the cost. When the cost value lies within the range $l-h^{-}\leq V_{c}(\rho)^{\pi}\leq l+h^{+}$, we apply gradient manipulation as defined in Eq. (12) to resolve the conflict between the two optimization objectives stated in Eq. (20) and Eq. (21).
A critical step in our algorithm is identifying safety violations, which involves determining whether $V_{c}^{\pi}$ exceeds the cost budget $l$. In the offline RL setting, off-policy collected datasets can lead to an underestimation of the cost value. Ideally, the estimated cost critic $\hat{Q}_{c}^{\pi}$ satisfies the asymptotic property $Q_{c}^{\pi}=\hat{Q}_{c}^{\pi}+\epsilon$, where $\epsilon$ is a zero-mean error term. However, under the expectation $\mathbb{E}_{s,a\sim\pi}[\min\hat{Q}_{c}^{\pi}(s,a)]\leq\min\mathbb{E}_{s,a\sim\pi}[Q_{c}^{\pi}(s,a)]$, the error term no longer has zero mean, resulting in an underestimation of the cost value Thrun and Schwartz (2014). This issue can be exacerbated by the temporal difference method, leading to an accumulated large bias. A common approach to address this challenge is to employ conservative Q-learning methods, which mitigate the underestimation issue at the cost of learning sub-optimal policies. The core difference between the reward critic and the cost critic is that the reward critic does not suffer from approximation errors as long as the relative priority between different actions is preserved. In contrast, underestimation of the cost critic may lead to unsafe policies. Therefore, we adopt a more conservative approach to evaluate safety conditions. Similar to the UCB (Upper Confidence Bound) method in bandit settings Hao et al. (2019), we use the UCB of the cost critic to assess safety during policy extraction. Specifically, we train an ensemble of cost critics $Q_{c}^{\pi,i}$, $i=1,2,\ldots,E$, and define the UCB cost critic as
$Q_{c}^{\pi,\text{UCB}}(s,a)=\bar{Q}_{c}^{\pi}(s,a)+k\cdot\operatorname{Std}_{i\in[E]}(Q^{\pi,i}_{c}(s,a)),$ (24)
where $\bar{Q}_{c}^{\pi}(s,a)$ represents the mean cost critic over the ensemble, $\operatorname{Std}_{i\in[E]}(Q^{\pi,i}_{c}(s,a))$ is the standard devistion of the $E$ ensembled critic estimators, and $k$ is a scaling factor controlling the confidence level. Using this substitution for the original cost critic, we can further ensure the implementation of safety at the expense of narrowing down the feasible policy set. Specifically, we define
$Q^{\pi,\text{UCB}}_{c}(\rho)=\mathbb{E}_{s\sim\rho,a\sim\pi}[Q^{\pi,\text{UCB}}_{c}(s,a)],$ (25)
and we compare it with the cost budget to determine the optimization objective. The implementation procedure is detailed in Algorithm 2 and Algorithm 3. Combined with the pretraining stage, our complete implementation is presented in Algorithm 1.
@@ 5.1 Performance On DSRL Benchmarks
Environments. We evaluate our method on the offline safe RL benchmark DSRL Liu et al. (2023a). We conduct extensive experiments on Safety-Gymnasium Ray et al. (2019) and Bullet-Safety-Gym Gronauer (2022). We evaluate the score of different methods using the normalized returns and normalized costs. The normalized returns and costs are defined as
$R_{\text{normalized}}=\frac{R_{\pi}-R_{\min}}{R_{\max}-R_{\min}},C_{\text{normalized}}=\frac{C_{\pi}-C_{\min}}{l+\epsilon}.$ (26)
@@
Here, $\epsilon$ is a regularizer for the case when $l=0$, and we set $\epsilon=0.1$. The reward $R_{\pi}$ is the accumulated return collected within an episode $R_{\pi}=\sum_{t=1}^{T}r_{t}$. Similarly, $C_{\pi}=\sum_{t=1}^{T}c_{t}$ is the accumulated cost collected within an episode. $R_{\max}$, $R_{\min}$ are the maximum and minimum accumulated returns of the offline dataset within an episode. We normalize the accumulated cost and return to better analyze the results. If the normalized cost is below 1.0, we can consider this policy as safe.
@@
Baseline Algorithms. We compare the performance of our algorithm against existing offline safe reinforcement learning (RL) algorithms under the soft constraints setting. The following six baseline algorithms are considered: 1) BC (Behavioral Cloning): Behavioral cloning imitates the behavior of the entire dataset to serve as a baseline. Variants include imitation of high-reward trajectories or safe trajectories. 2) BCQ-Lag Fujimoto et al. (2019): This approach extends behavioral cloning with Q-learning, framing safe reinforcement learning as a constrained optimization problem. A Lagrangian multiplier Stooke et al. (2020) is used to balance the reward maximization objective with the cost constraints. 3) BEAR-Lag Kumar et al. (2019): An extension of BEAR that incorporates a Lagrangian multiplier Stooke et al. (2020) to control cost constraints, enabling safe policy learning. 4) CDT (Constrained Decision Transformer) Liu et al. (2023b): An adaptation of the Decision Transformer Chen et al. (2021) for offline safe reinforcement learning, incorporating cost information into tokens to learn a safe policy. 5) CPQ (Conservative Policy Q-Learning) Xu et al. (2022): This method uses conservative Q-learning to pessimistically estimate the cost critic and updates the reward critic only with safe cost values, ensuring the policy adheres to safety constraints. 6) COptiDICE Lee et al. (2022): Based on the DICE algorithm, this method learns a stationary distribution for the safe reinforcement learning problem and extracts the optimal policy from this distribution.
Table 1: Training results across different tasks and algorithms. Safe policies are highlighted in blue, while unsafe policies are shown in gray. The table presents the normalized return and normalized cost for each task under different algorithms.
Reward (Cost) BC BCQ-Lag BEAR-Lag CDT CPQ COptiDICE Ours
AntVel 0.99(4.36) 0.99(8.40) -1.01(0.00) 0.91(6.04) -1.01(0.00) 0.992(11.41) 0.88(0.89)
HalfCheetahVel 0.976(19.04) 1.06(63.94) 0.947(104.45) 0.48(3.65) 0.71(14.7) 0.61(0.00) 0.86(0.00)
HopperVel 0.76(12.14) 0.81(14.92) 0.257(124.05) 0.21(1.12) 0.57(0.00) 0.14(7.84) 0.68(0.79)
Walker2DVel 0.78(5.38) 0.80(0.07) 0.763(0.8) 0.73(1.95) 0.08(0.96) 0.12(2.34) 0.74(0.30)
We present the comparison across different baselines in Mujoco-Gym environments. In Table 1, we present the normalized accumulated reward and cost per episode for each task. Overall, our algorithm demonstrates consistent safety across different tasks while achieving high rewards. Notably, in the Walker2D task, our method achieves a reward of 0.74, which is slightly below the 0.80 achieved by BCQ-Lagrangian. However, BCQ-Lagrangian fails to consistently ensure safe implementation in other tasks. In comparison, CPQ also achieves safety in most tasks but does so at the expense of reward due to its highly conservative nature, resulting in suboptimal policies in many cases. For CDT, we use the default parameter settings in the benchmark Liu et al. (2024) without fine-tuning the transformer’s hyperparameters. Overall, our method performs the best among all algorithms while reliably ensuring safe implementation.
5.2 Computational Efficiency
@@
We benchmark inference speeds across algorithms using 1,000 sample inputs on the HalfCheetah task. Apart from the 6 baselines above, We also compare the computation efficiency against constrained Diffuser Lin et al. (2023). While baseline methods like BCQ and CPQ achieve the fastest inference due to their lightweight MLP-based actors, they compromise safety or reward performance. In contrast, our algorithm strictly adheres to cost constraints without sacrificing reward quality. Compared to diffusion or transformer-based approaches our method are superior in inference speed, narrowing the gap between expressive generative models and efficient MLP architectures. The result is demonstrated in Figure 2.
Figure 2: Computational Efficiency vs. Performance Trade-off. Normalized score ($y$-axis, combining reward and safety metrics) versus inference time for generating 1,000 actions ($x$-axis).
5.3 Ablation Studies

### v2

@@ 4 Practical Algorithm
In this section, we present the detailed procedure for the implementation of our main algorithm Diffusion-Regularized Constrained Offline Safe Reinforcement Learning (DRCORL), as outlined in Algorithm 1, where we provide the SafeAdaptation step outlined in Algorithm 2 given in Appendix B. This includes both the pretraining stage and the policy extraction stage.
@@ 4.1 Pretraining the Diffusion Policy and Critics
$\min_{\pi\in\Pi}\mathbb{E}_{{s},{a},{s},a^{\prime}\sim\pi(\cdot|s^{\prime})}\big{[}\big{\|}c({s},{a})+\gamma Q^{\pi}_{c}(s^{\prime},a^{\prime})-Q^{\pi}_{c}(s,a)\|^{2}\big{]}-\alpha\mathbb{E}_{s,a\sim\pi}[Q^{\pi}_{c}(s,a)].$ (18)
@@ 4.2 Extracting Policy
Using the result obtained in Section 3.2, we can obtain the gradient of Eq. (19) and Eq. (20) with respect to $\theta$ using the score function of the diffusion model $\epsilon_{\psi}$ and the critic function. We take the Gaussian policy family under a constant variance. For example, with $\Pi:=\{\pi_{\theta}(a|s)=\mathcal{N}(a;m_{\theta}(s),\Sigma_{\theta}(s)\}$, we can simplify the gradient to
$g_{r}$ $=\mathbb{E}_{s,a\sim\pi_{\theta}}\left[\left(\nabla_{a}Q_{r}^{\pi_{\theta}}(s,a)+\tfrac{1}{\beta}h_{\psi}(s,a)\right)\nabla_{\theta}\pi_{\theta}(s)\right],$ (21)
$g_{c}$ $=\mathbb{E}_{s,a\sim\pi_{\theta}}\left[\left(-\nabla_{a}Q_{c}^{\pi_{\theta}}(s,a)+\tfrac{1}{\beta}h_{\psi}(s,a)\right)\nabla_{\theta}\pi_{\theta}(s)\right].$
where $h_{\psi}(s,a)=\frac{1}{\sqrt{\bar{\beta}_{t}}}\epsilon_{\psi}(a_{t},t|s)|_{t\rightarrow 0}$. In Eq. (21), $a_{t}$ denotes the action with perturbed noise, $a_{t}=\sqrt{\bar{\alpha}_{t}}a+\sqrt{\bar{\beta}_{t}}\epsilon_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,I)$, and the notation $\nabla_{\theta}\pi_{\theta}(s)$ denotes the gradient of the action given state $s$ with respect to $\theta$. By assuming $a=m_{\theta}(s)+\Sigma^{1/2}_{\theta}(s)z$ where $z\sim\mathcal{N}(0,I)$, we obtain that $\nabla_{\theta}\pi_{\theta}(s)=\nabla_{\theta}m_{\theta}(s)+\nabla_{\theta}\Sigma^{1/2}_{\theta}(s)z$. Our problem involves two competing objectives: maximizing reward (Eq. 10) and minimizing cost under safety violations (Eq. 20). To reconcile these goals, we adopt the gradient manipulation method from Gu et al. (2024a), originally proposed for online safe RL, as detailed in Algorithm 2. This method introduces slack variables $h^{-}$ and $h^{+}$. When $V_{c}^{\pi}(\rho)\leq l-h^{-}$, the policy is considered safe, and reward maximization is prioritized. If $V_{c}^{\pi}(\rho)\geq l+h^{+}$, we instead focus on cost minimization. Within the transition band $l-h^{-}\leq V_{c}^{\pi}(\rho)\leq l+h^{+}$, we apply gradient manipulation (Eq. 12) to balance the two objectives (Eqs. 19 and 20).
A key challenge is accurate safety assessment in the offline setting, where off-policy data may lead to cost underestimation. Ideally, the estimated critic satisfies $Q_{c}^{\pi}=\hat{Q}_{c}^{\pi}+\epsilon$, with $\epsilon$ zero-mean. However, under $\mathbb{E}_{s,a\sim\pi}[\min\hat{Q}_{c}^{\pi}(s,a)]\leq\min\mathbb{E}_{s,a\sim\pi}[Q_{c}^{\pi}(s,a)]$, the error becomes biased, leading to underestimated cost values Thrun and Schwartz (2014). Temporal difference learning can amplify this bias. While conservative Q-learning mitigates underestimation, it often yields sub-optimal policies. Crucially, unlike reward critics, where relative ranking suffices, underestimating cost critics can produce unsafe policies. To address this, we adopt a conservative evaluation approach inspired by the UCB (Upper Confidence Bound) technique Hao et al. (2019). Specifically, we train an ensemble of cost critics $Q_{c}^{\pi,i}$ for $i=1,\ldots,E$ and define the UCB estimator as $Q_{c}^{\pi,\text{UCB}}(s,a)=\bar{Q}_{c}^{\pi}(s,a)+k\cdot\operatorname{Std}_{i\in[E]}(Q^{\pi,i}_{c}(s,a)),$ where $\bar{Q}_{c}^{\pi}(s,a)$ is the ensemble mean and $k$ controls the confidence level. We then compute $Q^{\pi,\text{UCB}}_{c}(\rho)=\mathbb{E}_{s\sim\rho,a\sim\pi}[Q^{\pi,\text{UCB}}_{c}(s,a)]$ and compare it against the cost budget to determine whether to prioritize reward or cost. This full procedure is outlined in Algorithm 2.
@@ 5.1 Performance On DSRL Benchmarks
Environments. We evaluate our method on the offline safe RL benchmark DSRL Liu et al. (2023a). We conduct extensive experiments on Safety-Gymnasium Ray et al. (2019) and Bullet-Safety-Gym Gronauer (2022). We evaluate the score of different methods using the normalized returns and normalized costs. The normalized returns and costs are defined as $R_{\text{normalized}}=(R_{\pi}-R_{\min})/(R_{\max}-R_{\min}),C_{\text{normalized}}=(C_{\pi}-C_{\min})/(l+\epsilon)$. $\epsilon$ is a regularizer for the case when $l=0$, and we set $\epsilon=0.1$. The reward $R_{\pi}$ is the accumulated return collected within an episode $R_{\pi}=\sum_{t=1}^{T}r_{t}$. Similarly, $C_{\pi}=\sum_{t=1}^{T}c_{t}$ is the accumulated cost collected within an episode. $R_{\max}$, $R_{\min}$ are the maximum and minimum accumulated returns of the offline dataset within an episode. We normalize the accumulated cost and return to better analyze the results. If the normalized cost is below 1.0, we can consider this policy as safe.
@@
Baseline Algorithms. We compare the performance of our algorithm against existing offline safe reinforcement learning (RL) algorithms under the soft constraints setting. The following six baseline algorithms are considered: 1) BC-Safe (Behavioral Cloning): cloning the safe trajectories within the offline dataset. 2) BCQ-Lag Fujimoto et al. (2019): this approach extends behavioral cloning with Q-learning, framing safe reinforcement learning as a constrained optimization problem. A Lagrangian multiplier Stooke et al. (2020) is used to balance the reward maximization objective with the cost constraints. 3) BEARL Kumar et al. (2019): an extension of BEAR that incorporates a Lagrangian multiplier Stooke et al. (2020) to control cost constraints, enabling safe policy learning. 4) CDT (Constrained Decision Transformer) Liu et al. (2023b): an adaptation of the Decision Transformer Chen et al. (2021) for offline safe reinforcement learning, incorporating cost information into tokens to learn a safe policy. 5) CPQ (Conservative Policy Q-Learning) Xu et al. (2022): this method uses conservative Q-learning to pessimistically estimate the cost critic and updates the reward critic only with safe cost values, ensuring the policy adheres to safety constraints. 6) COptiDICE Lee et al. (2022): based on the DICE algorithm, this method learns a stationary distribution for the safe reinforcement learning problem and extracts the optimal policy from this distribution. 7) CAPS Chemingui et al. (2025): optimizing different constraints with shared representations. 8) CCAC Guo et al. (2025): safety RL algorithm matching state-action distributions and safety constraints. We present the comparison across different baselines in SafetyGym and BulletGym environments.
Result Analysis. Table 1 summarizes the normalized accumulated reward and cost per episode across tasks. Overall, our algorithm consistently achieves high rewards while reliably maintaining safety constraints. Notably, our method significantly outperforms baselines in tasks such as CarGoal1, CarGoal2, PointGoal1, PointGoal2, and BallCircle, ensuring both optimal reward and constraint satisfaction. However, it is fair to acknowledge that BCQ-Lag shows slightly superior reward performance in the Walker2dVel task, though such performance does not generalize reliably to other tasks, often violating safety constraints. Similarly, CDT and CAPS achieve good safety in some scenarios, such as CarGoal2 and AntVel, yet underperform significantly in reward optimization in other tasks. Overall, our method demonstrates a robust balance between safety and reward optimization across diverse benchmarks, highlighting its practical applicability.
Table 1: Normalized return ($\uparrow$: higher is better) and cost ($\downarrow$: lower is better; threshold at 1) for each policy across tasks. Results are averaged over three cost limit scales, 20 evaluation episodes and 5 random seeds. Policies with normalized cost $\leq 1$ (safe) are bolded; among these, the highest-reward safe policy per task is highlighted in blue. Policies with cost >1>1 (unsafe) are shown in gray.
BC-Safe BEARL BCQ-Lag CPQ COptiDICE CDT CAPS CCAC Ours
Task reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$ reward $\uparrow$ cost $\downarrow$
CarGoal1 0.38 0.46 0.71 4.45 0.46 3.27 0.68 3.73 0.48 2.31 0.65 3.75 0.40 1.35 0.84 6.52 0.91 0.00
CarGoal2 0.25 0.82 0.46 10.98 0.29 3.46 0.27 28.24 0.18 2.28 0.03 0.00 0.12 2.20 0.96 6.97 0.79 0.60
PointButton1 0.17 1.37 0.35 6.71 0.17 4.00 0.56 11.63 0.09 3.55 0.62 13.05 0.16 3.66 0.71 4.27 0.81 0.40
PointCircle1 0.88 3.79 -0.33 17.84 0.45 8.13 0.23 6.77 0.85 18.30 0.51 0.00 0.50 0.14 0.62 7.58 0.53 0.40
PointGoal1 0.53 0.88 0.76 3.46 0.59 5.06 0.41 0.69 0.52 5.26 0.68 4.35 0.20 0.53 0.77 5.20 0.88 0.00
PointGoal2 0.60 3.15 0.80 12.33 0.65 10.80 0.34 5.10 0.38 4.62 0.32 1.45 0.23 1.91 0.80 4.88 0.82 0.00
AntVel 0.87 0.52 -1.01 0.00 0.99 8.39 -1.01 0.00 0.99 11.41 0.91 0.97 0.81 0.36 0.71 0.39 0.88 0.89
HalfCheetahVel 0.94 0.71 0.95 104.45 1.06 63.94 0.71 14.70 0.61 0.00 0.96 0.61 0.88 0.22 0.85 0.93 0.86 0.00
HopperVel 0.21 1.50 0.26 124.05 0.81 14.91 0.57 0.00 0.14 7.83 0.21 1.12 0.83 0.00 0.11 0.68 0.68 0.79
Walker2dVel 0.78 0.08 0.76 0.80 0.80 0.07 0.08 0.95 0.12 2.34 0.73 1.95 0.78 0.00 0.21 0.26 0.74 0.30
CarRun 0.97 0.10 0.49 7.43 0.95 0.00 0.92 0.05 0.93 0.00 0.99 1.10 0.98 0.86 0.93 0.05 0.99 0.30
BallCircle 0.55 0.93 0.90 5.79 0.68 3.79 0.64 0.07 0.47 4.71 0.68 2.05 0.56 0.18 0.73 0.14 0.78 0.00
CarCircle 0.64 2.97 0.73 1.41 0.62 2.88 0.70 0.00 0.47 5.81 0.73 2.24 0.57 0.00 0.72 0.22 0.68 0.79
@@
Computational Efficiency. We benchmark inference speeds across algorithms using 1,000 sample inputs on the HalfCheetah task. Apart from the 6 baselines above, We also compare the computation efficiency against diffusion model-based algorithms TREBI Lin et al. (2023) and FISOR Zheng et al. (2024). While baseline methods like BCQ and CPQ achieve the fastest inference due to their lightweight MLP-based actors, they compromise safety or reward performance. In contrast, our algorithm strictly adheres to cost constraints without sacrificing reward quality. Compared to diffusion or transformer-based approaches our method are superior in inference speed, narrowing the gap between expressive generative models and efficient MLP architectures. The result is in Figure 2 (a).
5.2 Ablation Study